OpenStudy (anonymous):

Lee had a total of 40 e-mails on Monday. Eight of the e-mails were from her friend Cara. What percent of the e-mails were from Cara?

OpenStudy (mertsj):

8 out of 40? What percent is that?

OpenStudy (mertsj):

\[\frac{8}{40}=\frac{1}{5}=\frac{?}{100}\]

OpenStudy (anonymous):

what would the answer be?

OpenStudy (mertsj):

What do you get in place of the question mark?

OpenStudy (anonymous):

10???

OpenStudy (mertsj):

Can you reduce fractions?

OpenStudy (anonymous):

im so bad in math

OpenStudy (mertsj):

Can you reduce fractions?

OpenStudy (anonymous):

no

OpenStudy (mertsj):



OpenStudy (mertsj):

What should I multiply that 5 by so that I can get 100 for my new denominator?

OpenStudy (anonymous):

whats 100 divided by 5 ?

OpenStudy (anonymous):

whatever you find for that you multiply by one giving you your answer.

OpenStudy (anonymous):

would it be 20%?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ah thank you

OpenStudy (mertsj):

Ok. Amy and lg want to help you now. Too many cooks spoil the broth. Good Bye